# Spatially-Aware Diffusion Models with Cross-Attention for Global Field Reconstruction with Sparse Observations

### v1

@@ Abstract
Diffusion models have gained attention for their ability to represent complex distributions and incorporate uncertainty, making them ideal for robust predictions in the presence of noisy or incomplete data. In this study, we develop and enhance score-based diffusion models in field reconstruction tasks, where the goal is to estimate complete spatial fields from partial observations. We introduce a condition encoding approach to construct a tractable mapping mapping between observed and unobserved regions using a learnable integration of sparse observations and interpolated fields as an inductive bias. With refined sensing representations and an unraveled temporal dimension, our method can handle arbitrary moving sensors and effectively reconstruct fields. Furthermore, we conduct a comprehensive benchmark of our approach against a deterministic interpolation-based method across various static and time-dependent PDEs. Our study attempts to addresses the gap in strong baselines for evaluating performance across varying sampling hyperparameters, noise levels, and conditioning methods. Our results show that diffusion models with cross-attention and the proposed conditional encoding generally outperform other methods under noisy conditions, although the deterministic method excels with noiseless data. Additionally, both the diffusion models and the deterministic method surpass the numerical approach in accuracy and computational cost for the steady problem. We also demonstrate the ability of the model to capture possible reconstructions and improve the accuracy of fused results in covariance-based correction tasks using ensemble sampling.
@@ 1 Introduction
There are various ways to perform field reconstruction tasks with diffusion models that condition on observations. In the image processing domain, guided sampling or inpainting is frequently used because these techniques can be directly applied to trained diffusion models [36, 37, 33]. Guided sampling works by using the diffusion model to reconstruct unknown regions in the field. When applied in the physical domain, guided sampling is often combined with physical information to achieve physically realistic results [29, 30, 38, 32]. Some studies have also adopted the CFG method to incorporate sensing information [39, 40] or physical information [35, 29] as guiding information into diffusion models. This guiding information is typically integrated by augmenting the noise scale embedding and the latent representations of the fields. The cross-attention method is another frequently used approach for conditioning in the image processing domain [41, 42]. Cross-attention is a variant of self-attention [43] where the attention mechanism is applied between the image latent and the conditioning embedding. Compared to the augmentation performed in CFG, cross-attention has been shown to handle complex conditioning information [44], which could help capture variations in observation positions. Santos et al. [21] have investigated applying cross-attention-based deterministic method for field reconstruction tasks, and demonstrated promising results. However, to the best of our knowledge, cross-attention has not yet been explored in diffusion models for field reconstruction tasks.
@@
In this study, we propose a conditional encoding approach that leverages inductive bias and observation positions to construct a tractable mapping between observed and unobserved regions in a full-field diffusion model for field reconstruction tasks. We conduct an extensive benchmark comparing the diffusion model with the interpolation-based deterministic model, VCNN [1]. Furthermore, we evaluate different conditioning methods, including guided sampling (or inpainting), classifier-free guidance (CFG) [45], and cross-attention [46], while analyzing the effects of sampling hyperparameters. Our results indicate that applying cross-attention in conjunction with our proposed condition encoding block results in superior performance compared to the other two conditioning methods.
@@ 2.1 Problem formulation
Lef $\{s_{c,1},s_{c,2},\ldots,s_{c,N_{obs}}\}\subseteq\Omega$ denote the set of Voronoi-tessellated field of for the variable $\bm{x}_{c}\in\mathbb{R}^{N_{d}\times N_{d}}$. Each sub-region $s_{c,i}$ is defined as:
@@ 2.2 Diffusion model with spatial feature cross attention
Each model is trained for 100,000 steps on 8 Nvidia H100 GPUs, with weights updated using an Exponential Moving Average (EMA). We do not include a validation step for saving the best weights. Additional details on the training and implementation are provided in A.1.
@@ 3 Numerical Results
The reconstructed fields from noiseless observations using different methods are compared in Table 2. The results for the diffusion models are generated from an ensemble of 25 trajectories using the predictor-corrector scheme with 20 steps. We found that the predictor-corrector scheme generally provides more robust reconstructions than the multistep solver; reconstructions from the multistep solver are included in  C.1. Among the different conditioning methods, cross-attention consistently shows the best performance in terms of RMSE, nRMSE, and cRMSE across all PDEs.
@@
However, VT-UNet is more sensitive to observation noise levels, as shown in Figure  4. The diffusion models demonstrate more stable performance across different noise levels. When noise levels increase to $5\%$, the cross-attention method outperforms VT-UNet in all cases except for the CFD problem. The cross-attention method generally outperforms other conditioning methods across all observation noise levels. In contrast, the guided sampling method shows the worst performance for all PDEs and observation noise levels, indicating that for complex physical systems with sparse observations, guided sampling is insufficient to steer the sampling trajectory toward the correct solution. If unconditional generation capability is also desired, one should consider using the ControlNet approach [29] or setting the guidance scale in Equation (13) to be less than one.
@@
In the low-frequency stage, both the VT-UNet and the diffusion models effectively capture the underlying structure of the fields, with the difference between a single trajectory and the ensemble mean of the reconstructed fields being less significant. It is also noted that when the fields are sampled using the multistep solver, the diffusion models lose the ability to capture possible realizations, as shown in Figure 7. A comparison between different conditioning methods using multistep sampling is provided in C.1. The errors associated with multistep sampling are generally higher than those of the predictor-corrector method. Therefore, if a mean representation is desired, using the multistep solver can reduce computational cost with only a slight increase in error.
@@
Figure 9 displays the assimilated velocity field in the shallow water application following the diffusion model reconstruction. As mentioned in Section 2.3, the background error covariance in this case is empirically estimated from the ensemble generated using the diffusion model introduced in this paper. The ensemble size is fixed at 10 for all DA experiments. As can be clearly observed in Figure 9, the field reconstruction error is significantly reduced posterior to the DA process, particularly around the observable points. The estimated variance (i.e., the diagonal of the covariance matrix estimated using the 10 realizations of the diffusion model) is also shown in Figure 9.
@@ A.1 Implementation and Training details
The U-Net architecture is utilized for the denoiser function, $D_{\theta}$, following the same design as in [53]. The implementation is based on the PyTorch 2.3.1 and diffusers 0.29.2 libraries. The network consists of $N_{block}$ down-sampling and up-sampling blocks, where $N_{block}$ is the number of blocks required to achieve a bottleneck size of $8\times 8$. Specifically, $N_{block}$ equals 3 for the shallow water equations and 4 for the other problems. The first down-sampling block has 128 channels, while the remaining blocks have 256 channels. The up-sampling blocks are symmetric to the down-sampling blocks. The network uses a 3x3 convolutional layer with a stride of 2 for down-sampling and a nearest interpolation followed by a 3x3 convolutional layer for up-sampling. FiLM [57] is applied for both the noise level embedding and the classifier-free guidance (CFG).
@@ C.1 Diffusion model: Multistep sampling
The results for the diffusion models with the multistep sampling scheme are shown in Table 4. These results are generated from an ensemble of 25 trajectories using the predictor-corrector scheme with 20 steps. In general, the cross-attention method demonstrates the best performance in terms of RMSE, nRMSE, and cRMSE across all PDEs.

### v2

@@ Abstract
Diffusion models have gained attention for their ability to represent complex distributions and incorporate uncertainty, making them ideal for robust predictions in the presence of noisy or incomplete data. In this study, we develop and enhance score-based diffusion models in field reconstruction tasks, where the goal is to estimate complete spatial fields from partial observations. We introduce a condition encoding approach to construct a tractable mapping between observed and unobserved regions using a learnable integration of sparse observations and interpolated fields as an inductive bias. With refined sensing representations and an unraveled temporal dimension, our method can handle arbitrary moving sensors and effectively reconstruct fields. Furthermore, we conduct a comprehensive benchmark of our approach against a deterministic interpolation-based method across various static and time-dependent PDEs. Our study attempts to addresses the gap in strong baselines for evaluating performance across varying sampling hyperparameters, noise levels, and conditioning methods. Our results show that diffusion models with cross-attention and the proposed conditional encoding generally outperform other methods under noisy conditions, although the deterministic method excels with noiseless data. Additionally, both the diffusion models and the deterministic method surpass the numerical approach in accuracy and computational cost for the steady problem. We also demonstrate the ability of the model to capture possible reconstructions and improve the accuracy of fused results in covariance-based correction tasks using ensemble sampling.
@@ 1 Introduction
There are various ways to perform field reconstruction tasks with diffusion models that condition on observations. In the image processing domain, guided sampling or inpainting is frequently used because these techniques can be directly applied to trained diffusion models [36, 37, 33]. Guided sampling works by using the diffusion model to reconstruct unknown regions in the field. When applied in the physical domain, guided sampling is often combined with physical information to achieve physically realistic results [29, 30, 38, 32]. Some studies have also adopted the classifier-free guidance (CFG) [39] method to incorporate sensing information [40, 41] or physical information [35, 29] as guiding information into diffusion models. This guiding information is typically integrated by augmenting the noise scale embedding and the latent representations of the fields. The cross-attention method is another frequently used approach for conditioning in the image processing domain [42, 43]. Cross-attention is a variant of self-attention [44] where the attention mechanism is applied between the image latent and the conditioning embedding. Compared to the augmentation performed in CFG, cross-attention has been shown to handle complex conditioning information [45], which could help capture variations in observation positions. Santos et al. [21] have investigated applying cross-attention-based deterministic method for field reconstruction tasks, and demonstrated promising results. However, to the best of our knowledge, cross-attention has not yet been explored in diffusion models for field reconstruction tasks.
@@
In this study, we propose a conditional encoding approach that leverages inductive bias and observation positions to construct a tractable mapping between observed and unobserved regions in a full-field diffusion model for field reconstruction tasks. We conduct an extensive benchmark comparing the diffusion model with the interpolation-based deterministic model, VCNN [1]. Furthermore, we evaluate different conditioning methods, including guided sampling (or inpainting), CFG [39], and cross-attention [46], while analyzing the effects of sampling hyperparameters. Our results indicate that applying cross-attention in conjunction with our proposed condition encoding block results in superior performance compared to the other two conditioning methods.
Field reconstruction from sparse sensor data is closely related to super-resolution tasks, where the goal is to recover high-resolution fields from limited observations, as commonly seen in fluid dynamics applications [47]. Recent studies have applied diffusion models to super-resolution problems, demonstrating their ability to handle complex nonlinear structures [35, 38]. Although our work focuses on field reconstruction, the proposed cross-attention method can also be applied to super-resolution tasks, given the similarity in input data structure and the need for high-fidelity reconstruction from sparse observations.
@@ 2.1 Problem formulation
Let $\{s_{c,1},s_{c,2},\ldots,s_{c,N_{obs}}\}\subseteq\Omega$ denote the set of Voronoi-tessellated field for the variable $\bm{x}_{c}\in\mathbb{R}^{N_{d}\times N_{d}}$. Each sub-region $s_{c,i}$ is defined as:
@@ 2.2 Diffusion model with spatial feature cross attention
Each model is trained for 100,000 steps on 8 Nvidia H100 GPUs, with weights updated using an Exponential Moving Average (EMA). We do not include a validation step for saving the best weights. Additional details on the training and implementation are provided in A.1. A discussion of how the trained models can be finetuned to handle different resolutions is provided in C.4.
@@ 3 Numerical Results
The reconstructed fields from noiseless observations using different methods are compared in Table 2. The results for the diffusion models are generated from an ensemble of 25 trajectories using the predictor-corrector scheme with 20 steps. We found that the predictor-corrector scheme generally provides more robust reconstructions than the multistep solver [62]; reconstructions from the multistep solver are included in  C.1. Among the different conditioning methods, cross-attention consistently shows the best performance in terms of RMSE, nRMSE, and cRMSE across all PDEs.
@@
However, VT-UNet is more sensitive to observation noise levels, as shown in Figure  4. The diffusion models demonstrate more stable performance across different noise levels. When noise levels increase to $5\%$, the cross-attention method outperforms VT-UNet in all cases except for the CFD problem. The cross-attention method generally outperforms other conditioning methods across all observation noise levels. In contrast, the guided sampling method shows the worst performance for all PDEs and observation noise levels, indicating that for complex physical systems with sparse observations, guided sampling is insufficient to steer the sampling trajectory toward the correct solution. A comparison of the diffusion-reaction problem under colored noise conditions is provided in C.5. If unconditional generation capability is also desired, one should consider using the ControlNet approach [29] or setting the guidance scale in Equation (13) to be less than one.
@@
In the low-frequency stage, both the VT-UNet and the diffusion models effectively capture the underlying structure of the fields, with the difference between a single trajectory and the ensemble mean of the reconstructed fields being less significant. It is also noted that when the fields are sampled using the multistep solver, the diffusion models lose the ability to capture possible realizations, as shown in Figure 7. A comparison between different conditioning methods using multistep sampling is provided in C.1. The errors associated with multistep [62] sampling are generally higher than those of the predictor-corrector method, although multistep sampling achieves better nRMSE for the fine-tuned model (see C.4). Therefore, if a mean representation is desired, using the multistep solver can reduce computational cost with only a slight increase in error.
@@
Figure 9 displays the assimilated velocity field in the shallow water application following the diffusion model reconstruction. As mentioned in Section 2.3, the background error covariance in this case is empirically estimated from the ensemble generated using the diffusion model introduced in this paper. The ensemble size is fixed at 10 for all DA experiments. As can be clearly observed in Figure 9, the field reconstruction error is significantly reduced posterior to the DA process, particularly around the observable points. The estimated variance (i.e., the diagonal of the covariance matrix estimated using the 10 realizations of the diffusion model) is also shown in Figure 9. An energy plot of the shallow water equation is provided in C.6, demonstrating that the models maintain consistent performance over time, even when each time step is treated as a separate instance.
@@ A.1 Implementation and Training details
The U-Net architecture is utilized for the denoiser function, $D_{\theta}$, following the same design as in [54]. The implementation is based on the PyTorch 2.3.1 and diffusers 0.29.2 libraries. The network consists of $N_{block}$ down-sampling and up-sampling blocks, where $N_{block}$ is the number of blocks required to achieve a bottleneck size of $16\times 16$, where $N_{block}$ is the number of blocks required to achieve a bottleneck size of $16\times 16$, the last block in the down-sampling part and the first block in the up-sampling part do not have down-sampling and up-sampling layers, respectively. Specifically, $N_{block}$ equals 3 for the shallow water equations and 4 for the other problems. The first down-sampling block has 128 channels, while the remaining blocks have 256 channels. The up-sampling blocks are symmetric to the down-sampling blocks. The network uses a 3x3 convolutional layer with a stride of 2 for down-sampling and a nearest interpolation followed by a 3x3 convolutional layer for up-sampling. FiLM [58] is applied for both the noise level embedding and the classifier-free guidance (CFG).
Zero padding in CNNs can leak positional information [63], and typically, special treatment is required for processing different types of periodic conditions [64]. In this work, we do not implement these techniques; instead, we rely on skip connections and self-attention mechanisms to capture global information. However, we expect that periodic padding could improve the generalizability of the models and could be explored in future work.
@@ C.1 Diffusion model: Multistep sampling
The results for the diffusion models with the multistep [62] sampling scheme are shown in Table 4. These results are generated from an ensemble of 25 trajectories using the predictor-corrector scheme with 20 steps. In general, the cross-attention method demonstrates the best performance in terms of RMSE, nRMSE, and cRMSE across all PDEs.
@@ C.2 Diffusion model: Generating not Memorizing
C.4 Varying resolution
CNN-based models have been shown to generalize well across different resolutions [69, 70], and the backbone of our diffusion model is UNet, which is an variant of CNN with skip connection at different hierarchical levels. However, in our investigation with the higher resolution darcy flow problem, we found that both VT-UNet and the conditional diffusion model fails to generalize. This is likely due to the changes in feature scale perceived by the convolutional kernels as the resolution increases. To address this issue, popular approaches include scaling input, training with multiple resolutions (data augmentation), and fine-tuning the model on the target resolution.
Figure 18: Step loss comparison between training from scratch and fine-tuning on the Darcy flow problem with resolution of $256\times 256$. The fine-tuning process shows a faster convergence rate. Step loss is smoothed with a moving average of 10 steps.
Table 5: Comparison of training from scratch and fine-tuning on the Darcy flow problem with resolution of $256\times 256$ on the 1k test data with an ensemble size of 25. The number of sampled points is 256, which is equivalent to the 1.37% in $128\times 128$ resolution.
Model Sampling RMSE nRMSE cRMSE
Scratch Predictor-Corrector 0.177 0.133 0.018
Scratch Multistep 0.321 0.234 0.075
Finetune Predictor-Corrector 0.249 0.187 0.019
Finetune Multistep $\bm{0.166}$ $\bm{0.123}$ $\bm{0.016}$
The $256\times 256$ resolution Darcy flow problem is generated with the same hyperparameters and seeding as the $128\times 128$ resolution. Both models are trained for 20,000 steps with a learning rate of $51e-$ 5. The smoothed step loss is shown in Figure 18. The fine-tuning process demonstrates a faster convergence rate. We compare the performance of training from scratch and fine-tuning from the $128\times 128$ resolution in Table 5. Interestingly, the model trained from scratch shows better performance with the predictor-corrector sampling method, which aligns with the findings in the main text. However, the trend reverses for the fine-tuned model, where the multistep method outperforms the predictor-corrector method. This indicates that the trained model can be adapted to different resolutions, and sampling methods should be revised accordingly.
C.5 Colored noise
Figure 19: Colored noise performance comparison for the diffusion reaction problem. Results of nRMSE are computed with 1k test data with an ensemble size of 25 and 1% known values.
Results of 1D diffusion-reaction problems with colored noise are shown in Figure 19. The performance of the diffusion model is evaluated with respect to noise levels, where the nRMSE is computed for 1k test data with an ensemble size of 25 and 1% known values. In general, the same tendency is observed across different types of noise, where the performance of the diffusion model surpasses the VT-UNet baseline as the noise level increases.
C.6 Shallow water energy plot
Figure 20: Energy plot of shallow water equations on 10 test set trajectories, with 1% values known, the cross-attention diffusion model is utilized with an ensemble size of 25 and 20 reverse sampling steps.
In Figure 20, we present the normalized energy plot for the shallow water equations on 10 test set trajectories, with 1% known values, 0% noise, and the cross-attention diffusion model employing an ensemble of 25 and 20 reverse sampling steps. This plot supports the findings in the main text, where nRMSE results show that the VT-UNet model performs better under noise-free conditions. The low RMSE observed in these noise-free scenarios aligns with this energy-based evaluation, highlighting the model’s effectiveness in capturing fine-scale features, even with sparse observations.